OpenStudy (anonymous):

what is the polynomial function in standered form with zeros 1,2,-2, and -3

OpenStudy (campbell_st):

well its a quartic so well is a zero is at x = 1 then x -1 is a factor x =2 x -2 is a fator etc P(x) = (x -1)(x -2)(x+2)(x +3) so just distribute for the answer

OpenStudy (anonymous):

I dont know how to distribute